I need some help trying to sovle this problom every time i try to solve it. it comes out to 5 but thats wrong tho.
4(a-2)=2(a-1)

Answer by pwac(253)

4(a-2)=2(a-1) first remove brackets by multiplying by number in front
4a-8=2a-2 take a terms to left and numbers to right
4a-2a=-2+8 remember sign changes when term changes side
2a=6 now divide both sides by 2
2a/2=6/2
a=3
hope you understand
Pete